OpenStudy (anonymous):

Please help! Perform the indicated operation. Leave all answers in the form a + b i. (1 - i ) (2 + 3 i ) a. -1 - 5 i b. -1 + i c. 5 + i d. 5 - 5 i

hartnn (hartnn):

do u know this: (a+b)(c+d)=ac+ad+bc+bd ?

OpenStudy (anonymous):

I'm familiar. Hmmm, so that's what I use?

hartnn (hartnn):

and also the fact that i*i = -1 just try it....

mathslover (mathslover):

\[\large{(1-i)(2+3i)=1(2+3i)-i(2+3i)}\] \[\large{(1-i)(2+3i)=1(2)+1(3i)-i(2)-i(3i)}\] this is the use of taht @IloveCharlie

mathslover (mathslover):

*that

OpenStudy (anonymous):

Hmm... :/

OpenStudy (anonymous):

d.

mathslover (mathslover):

what is i(3i) ?

mathslover (mathslover):

i * 3i = 3 i ^2 = 3(-1) = -3 2+3i-2i+3 5+i

OpenStudy (anonymous):

i(3i)=-3 @mathslover

OpenStudy (anonymous):

9+2i?

OpenStudy (anonymous):

:/

hartnn (hartnn):

@mathslover already gave u answer,see the comments....

OpenStudy (anonymous):

Yes but I can't match it up :/

hartnn (hartnn):

he gave u answer as 5+i,which is option c here are all steps together: (1−i)(2+3i) =1(2)+1(3i)−i(2)−i(3i) =2+3i-2i+3 =5+i

OpenStudy (anonymous):

Oh I understood I just must have missed it. Thank you for explaining the steps BOTH of you I appreciate it

mathslover (mathslover):

Oh sorry yes : i *3 i =3 i^2 = -3